I am having trouble solving these types of problems. Please help me understand by explaining and showing me the steps and methods needed to work this out. thank you.

I am having trouble solving these types of problem
Answer #1

R, il R, 12

Use kIrchoff's law in left loop:
-E1 + i2*(R1) + (i2-i1)*R2 = 0
-10 + i2*5 + (i2-i1)*5=0
-5* i1 + 10*i2 = 10 ....eqn 1


Use kIrchoff's law in lower loop:
-E2 + (i1-i2)*R2 + i1*R3 =0
-15 + (i1-i2)*5 + i1*5 =0
10i1 - 5i2 = 15 ...eqn 2

solving eqn 1 and eqn 2,
i1= 2.7 A
i2= 2.3 A

Current in left branch is i2
Answer: 2.3 A